OpenStudy (anonymous):

log 5^64-log 5^ (8/3)+ log5 ^ 2=log5^4p

OpenStudy (campbell_st):

the easiest way to solve this problem is to use the log laws for powers \[\log (x^a) = a \log (x)\] so you will have \[64 \log 5 - \frac{8}{3} \log 5 + 2 \log 5 = 4p \log 5\] divide log 5 as a common factor and then solve for p

OpenStudy (anonymous):

the answer i got for it was 12.

OpenStudy (campbell_st):

66 - 8/3 = 4p so p = 16 1/2 - 2/3 looks like 15 and a bit to me